# Parameter-Specific Bias Diagnostics in Random-Effects Panel Data Models

### v2

@@ Abstract.
The Hausman specification test detects inconsistency in random-effects estimators by comparing them with alternative fixed-effects estimators. This note shows how a recently proposed bias diagnostic for linear mixed models can complement this test in random-effects panel-data applications. The diagnostic delivers parameter-specific internal estimates of finite-sample bias, together with permutation-based $p$-values, from a single fitted random-effects model. We illustrate its use in a gasoline-demand panel and in a value-added model for teacher evaluation, using publicly available R packages, and we discuss how the resulting bias summaries can be incorporated into routine practice.
@@ 2.2. Bias Diagnostic
The random-effects (RE) estimator of $\boldsymbol{\beta}$ is
(2) $\hat{\boldsymbol{\beta}}_{\text{RE}}=({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{y}},$
with $\hat{{\mathbf{V}}}={\mathbf{Z}}\hat{{\mathbf{G}}}{\mathbf{Z}}^{\prime}+\hat{\mathbf{R}}$. Following Karl & Zimmerman (2021), we treat $\mathbf{X}$} as fixed and $\mathbf{Z}$} as stochastic. Let ${\mathbf{V}}={\mathbf{Z}}{\mathbf{G}}{\mathbf{Z}}^{\prime}+\mathbf{R}$ denote the true marginal covariance matrix of $\mathbf{y}$}, and define
$\boldsymbol{\nu_{k}}=\mathbf{k}^{\prime}({\mathbf{X}}^{\prime}{\mathbf{V}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}{\mathbf{V}}^{-1}{\mathbf{Z}}.$
Karl & Zimmerman (2021) show that for any $p\times 1$ vector $\mathbf{k}$ such that $\mathbf{k}^{\prime}\boldsymbol{\beta}$ is an estimable linear combination of the fixed effects (for example, a single regression coefficient $\beta_{j}$ or a contrast such as $\beta_{j}-\beta_{\ell}$),
(3) $\operatorname{Bias}(\mathbf{k}^{\prime}\hat{\boldsymbol{\beta}}_{\text{RE}})=\mathbf{E}[\boldsymbol{\nu_{k}}^{\prime}\boldsymbol{\eta}].$
Karl & Zimmerman (2021) propose a plug-in approximation based on the fitted mixed model. Writing
(4) $\hat{\boldsymbol{\nu_{k}}}=\mathbf{k}^{\prime}({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{Z}},$
and denoting the empirical best linear unbiased predictors (EBLUPs) of the random effects by $\hat{\boldsymbol{\eta}}$, the plug-in quantity $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$ serves as an internal estimate of the bias of $\mathbf{k}^{\prime}\hat{\boldsymbol{\beta}}_{\text{RE}}$. The exact bias is zero if any of the following three conditions hold: (i) ${\mathbf{X}}^{\prime}\mathbf{R}^{-1}{\mathbf{Z}}=\boldsymbol{0}$, (ii) the random-effects vector $\boldsymbol{\eta}$ is independent of $\mathbf{Z}$}, or (iii) $\boldsymbol{\eta}$ is treated as fixed (producing the FE estimator). The third condition implies that the FE estimator remains unbiased in settings where the RE estimator is biased, so that the internal bias estimate often approximates the difference between the RE and FE estimators (Karl & Zimmerman, 2021).
@@ 3. Gasoline Consumption Data Application
Next, we fit the mixed model with lme4 and run the bias diagnostic (Listing 2). For comparison of parameter estimates, we also fit the corresponding fixed effects model, although this is not required by the bias test.
@@
The graphical summaries of the bias estimates and reference distributions appear in Figure 1. The results from Listing 2 are combined into Table 1. The Hausman test indicates global inconsistency of the RE estimators, and the bias diagnostic suggests that the RE estimate of lrpmg is significantly biased downward. Note the similarity of the internal bias estimates from the RE model to the RE–FE differences in Table 1.
@@ 4. Value-added Model (VAM) Results with GPvam
The CP estimates and bias diagnostics for the fixed effects appear in Table 2. The Hispanic parameter estimate is significantly downward-biased, while the White and Asian/Pac Island parameter estimates are significantly upward-biased. The contrast White - Hispanic reveals even stronger bias with a point estimate of $0.1287$ and a $p$-value of 0 (none of the one million permutations exceeded the test statistic). Although the bias is statistically significant, its magnitude may not be practically significant for this particular application. Nevertheless, this example illustrates how the diagnostic can facilitate the investigation of potential bias due to nonrandom allocation of random effects on a case-by-case basis. In this example we do not propose a change in estimator based solely on the diagnostic. Instead, the bias estimates are used descriptively to flag coefficients and contrasts whose interpretation may be sensitive to nonrandom teacher assignment.

### v3

@@ Abstract.
The Hausman specification test detects inconsistency of the random-effects estimator by comparing it with an alternative fixed-effects estimator. This note shows how a recently proposed bias diagnostic for linear mixed models can complement this test in random-effects panel-data applications. The diagnostic delivers parameter-specific internal estimates of finite-sample bias of the random-effects estimator, together with permutation-based $p$-values, from a single fitted random-effects model. We illustrate its use in a gasoline-demand panel and in a value-added model for teacher evaluation, using publicly available R packages, and we discuss how the resulting bias summaries can be incorporated into routine practice.
@@ 2.2. Bias Diagnostic
Let $\hat{{\mathbf{G}}}$ and $\hat{\mathbf{R}}$ denote estimates of $\mathbf{G}$} and $\mathbf{R}$, and define $\hat{{\mathbf{V}}}={\mathbf{Z}}\hat{{\mathbf{G}}}{\mathbf{Z}}^{\prime}+\hat{\mathbf{R}}$. The empirical random-effects (RE) estimator of $\boldsymbol{\beta}$ is
(2) $\hat{\boldsymbol{\beta}}_{\text{RE}}=({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{y}}.$
Following Karl & Zimmerman (2021), we treat $\mathbf{X}$} as fixed and $\mathbf{Z}$} as stochastic. For any $p\times 1$ vector $\mathbf{k}$ such that $\mathbf{k}^{\prime}\boldsymbol{\beta}$ is an estimable linear combination of the fixed effects (for example, a single regression coefficient $\beta_{j}$ or a contrast such as $\beta_{j}-\beta_{\ell}$), define
(3) $\hat{\boldsymbol{\nu_{k}}}=\mathbf{k}^{\prime}({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{Z}}.$
Karl & Zimmerman (2021) show that, under regularity conditions,
(4) $\operatorname{Bias}(\mathbf{k}^{\prime}\hat{\boldsymbol{\beta}}_{\text{RE}})=\mathbf{E}[\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\eta}].$
The bias of the random-effects estimator $\mathbf{k}^{\prime}\hat{\boldsymbol{\beta}}_{\text{RE}}$ is zero if either (i) ${\mathbf{X}}^{\prime}\mathbf{R}^{-1}{\mathbf{Z}}=\boldsymbol{0}$ or (ii) the random-effects vector $\boldsymbol{\eta}$ is independent of $\mathbf{Z}$}. A third way to avoid bias in practice is to treat $\boldsymbol{\eta}$ as fixed (producing the FE estimator), in which case the corresponding fixed-effects estimator of $\mathbf{k}^{\prime}\boldsymbol{\beta}$ remains unbiased.
@@ 3. Gasoline Consumption Data Application
Next, we fit the mixed model with lme4 and run the bias diagnostic (Listing 2). For comparison of parameter estimates, we also fit the corresponding fixed effects model, although this is not required by the bias test. Although the random-effects estimator from plm and the random-intercept estimator from lmer are not exactly identical, they are very similar in this simple random-intercept setting. We therefore use plm for the Hausman test and lmer for the bias diagnostic, since mixedbiastest is implemented on top of lme4.
@@
The graphical summaries of the bias estimates and reference distributions appear in Figure 1. The results from Listing 2 are combined into Table 1. The Hausman test rejects the null hypothesis that the RE estimator is consistent, and the bias diagnostic suggests that the RE estimator for lrpmg is downward biased. Note the similarity of the internal bias estimates from the RE model to the RE–FE differences in Table 1.
@@ 4. Value-added Model (VAM) Results with GPvam
The CP estimates and bias diagnostics for the fixed effects appear in Table 2. The diagnostic indicates that the estimator for the Hispanic effect is downward biased, while the estimators for the White and Asian/Pac Island effects are upward biased. The diagnostic for the contrast White - Hispanic reveals even stronger bias: the estimated contrast is $0.1287$, and none of the one million permutations exceeded the test statistic ($p$-value $\approx 0$). Although the bias is statistically significant, its magnitude may not be practically significant for this particular application. Nevertheless, this example illustrates how the diagnostic can facilitate the investigation of potential bias due to nonrandom allocation of random effects on a case-by-case basis. In this example we do not propose a change in estimator based solely on the diagnostic. Instead, the bias estimates are used descriptively to flag coefficients and contrasts whose interpretation may be sensitive to nonrandom teacher assignment.